OpenStudy (anonymous):

A coin is tossed 11 times. What is the probability of obtaining exactly 9 heads? A. 0.0269 B. 0.0054 C. 0.1611 D. 0.0806

OpenStudy (kc_kennylau):

\[\Large P=\left(\begin{array}{c}11\\9\end{array}\right)\times\left(\frac12\right)^{11}\]

OpenStudy (anonymous):

0.0054 (B)

OpenStudy (kc_kennylau):

This is not what I have calculated

OpenStudy (anonymous):

Before i got A(0.0269), could that be correct?

OpenStudy (phi):

11C9 = 55 ½^11 is 1/2048

OpenStudy (anonymous):

?

OpenStudy (kc_kennylau):

Yep I got A

OpenStudy (anonymous):

thank you!

OpenStudy (kc_kennylau):

no problem